A garden table and a bench cost 681 combined. The garden table costs 69
less than the bench. What is the cost of the bench?

Answer by ikleyn(52835)

.
A garden table and a bench cost 681 combined. The garden table costs 69
less than the bench. What is the cost of the bench?
~~~~~~~~~~~~~~

Write equations as you read the problem

    B + T = 681     (1)   (total cost)

    B - T =  69     (2)   (T costs 69 less than B)


To solve, add equations.  The terms with "T" will annihilate, and you will get then

    2B = 681 + 69 = 750  --->  B = 750/2 = 375.


Now from equation (1)

         T = 681 - B = 681 - 375 = 306.


ANSWER.  Table costs 306;  B costs 375.

Solved.